OpenStudy (anonymous):

Solve for exact answers: 6(sqrtx)+x=3

OpenStudy (earthcitizen):

\[6\sqrt{x}+x=3\]

OpenStudy (earthcitizen):

\[6\sqrt{x}=3-x\] \[x=(3-x)^{2}/36\]

OpenStudy (earthcitizen):

\[x^{2}-6x-36x+9=0\]

OpenStudy (earthcitizen):

\[x ^{2}-40x+9=0\]

OpenStudy (anonymous):